FEL Upgrade Project Weekly Brief
May 14-18, 2001

Our current analysis (by Dave Douglas) of the energy acceptance of the electron beam transport system in the recirculator is looking very good compared to the nominal design value of 10%. See the WBS 3 report for details.

After the receipt of the project financial reports for the month of April, we began the process of compiling the monthly update to the cost-performance report.

G. Neil continues preparation of  material for his presentation on high power FEL scaling for the upcoming
June 5-6 workshop which is being sponsored by the JTO.

We have yet not received funds through our ONR-DOE contract for the JTO funded proposal on laser damage studies and have queried ONR for release of funds so that this work can proceed.

WBS 3 (Beam Physics):
Correction of a misperception in the DIMAD analysis of recirculator performance has suggested the driver may accept an FEL-induced momentum spread in excess of the 10% design specification. As a result, the system may allow operation at FEL extraction efficiencies well beyond the nominal 1% value, with an associated increase in output power. (Editorial note from the project manager: THIS IS A VERY IMPORTANT RESULT!-every once in a while Maxwell’s Demon takes a bat swing in the right direction.

Estimates characterizing pi-bend specifications were documented in JLAB-TN-01-024.

WBS 4 (Injector):
The gun drawings are back to the designer after corrections were made. The drawings are going to a manufacturing sequence series. Ran voltage on the 1um SST implanted electrode ( implanted with a faraday shield in place). Preliminary results show much higher emission currents than the electrodes implanted without the faraday shield. The anode plate is being polished to 1 um.

Gun HVPS – Glassman, Inc. expects to ship by the end of the month.

WBS 5 (SRF):
Miscellaneous cryogenic piping package complete and the SOW has been signed off.  Received and evaluating quotes for tuner fabrication.

WBS 6 (RF):
Zone 1 - Both HVPS's were cleaned inside of any dust and debris of prevent random arcing.  The brushes for unit 4 were replaced during the last run.  Any bad brushes in unit 3 will be changed next week.

Zone 2 - Operational.

Zone 3 - No progress.  Waiting for additional parts.

Zone 4 - Except for position #8, the zone is complete and waiting for a cryomodule.  This week, 7 of the klystrons were operated to determine the stability of their cathode current and hence RF power when the heater voltage is set much less than normal.  The results are not yet available.

100 kW Klystrons - CPI has not yet been able to determine if the vacuum problem is due to the copper or the test fixture.  The additional tests will be run next week.

CPS for 100 kW Klystrons - Continuing to contact transformer manufacturers.

WBS 8 (Instrumentation):
A productive visit was made to Triumf regarding their CAN bus implementation.  They have ~ 500 devices installed and operational on the Controller Area Network (CAN). It is our intention to use this for the new BPM electronics and for the power supply control for the upgrade.  The new VME CAN interface also arrived this week, it is the same one that is used at Triumf (and supported in EPICS).  All major components are on site for this BPM prototype, the '807 DSP (samples just released from production this week), the Philips CAN interface (to the DSP), the RF filters, and the Analog devices RF-to-DC converter. The only thing that remains is the 3/3.3 volt power management/conversion.  These will be tested during the summer run.

The Picomotor Mux Chassis has been re-installed in the control room.  Dan put together the shelves in Lab 1 to help clear out Lab 6. A video cable was installed from the switcher to FL07B02 for the SPIRICON.  James added the Alignment Mode Chassis and the User Lab Status Indicator Systems to the LSS System Block Diagram.  They also raised the legs on the HP workstation to support getting that up and running.  James put together a daughter board for
the OBPMs for testing and Dan has stuffed another 64 Bit interface card to use in conjunction  with the testing.

Mike and Patrick replaced all of the homemade VME panels with ones we had made.  Mike also has a data base inventory of all of the VME panels.  They are divided into: installed and spare with the spares divided into those that are stuffed, partially stuffed and blank.  Patrick has the rabbit board running a small motor and is expanding the project to include memory use. The front and rear panels for the user power supplies came back from the vendor and those are being assembled in the lab in the ARC.  I expect they will be completed either today or Monday.  The Lab in room 307 in the ARC was cleaned prior to assembly beginning.

The Machine Protection System drawing work continues.  Most of the rear of the panel has been documented and system interconnections are being labeled and cataloged as work progresses.  Several drawings are in EECAD:  Ion Pump Power Supply Main Board Schematic; Molectron OBPM and daughter board Schematic; 64 Bit General purpose Digital I/O card; Single channel OBPM Amplifier Schematic; Picomotor Relay and Ion Pump Power Supply Wiring Diagrams and the Analog Differential Driver Board Schematic as well as several pages of the MPS System Drawing.  The following drawings have been received for review/ signature: OBPM enclosure box; Upstairs Dump Raster Board Schematic; Picomotor Relay Chassis Mechanical; and the LSS System Block Diagram.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o DULY Research is making small corrections to their magnetic model of the small injector Dipole (GV) so that it
   matches the excitation of the large injector dipole without a shunt.  They are also doing sensitivity studies of the
   position of the field clamps on field integral for both magnets.
o  They continue work on the drawings of the DV. A practical hardware design.
o  They delivered the preliminary details and assembly of the large injector dipole (GU) for checking.
Optical Chicane Dipole (DW)
o  We received the electrical steel pole tip material from the custom specified roll that had the best thickness
    uniformity characteristic.
Arc Dipoles (GY, GX, GQ)
o  The magnetic analysis of the GY is priority, concentrating on trying small changes to electrical steel edge near the
    holes to minimize their effects.  Tom Schultheiss is will come to JLab to finalize the magnetic analysis, aiming for
    next week.
o  We checked about half of the 3/4 inch stack of drawings of the GY magnet.
o  They are continuing design and magnetic modeling the Arc Bend Dipole (GX), getting the model to work, but
    not getting the expected 1 part in 10000 quality.  Tom is making the mesh finer at the poles.  In the design area,
    the assembly drawing and a considerable number of the detail drawings took shape as preliminary drawings.
QG (3 inch quad)
o  The coil manufacturing package remains in Procurement cycle.
o  The core drawing and statement of work were signed as the procurement package was made ready to send out
o  The prototype quadrupole had two more chamfer cuts, winding up at .375 inch.  Extension of the results
    indicates the ideal is around .385 where the sum of the multipoles is around a minimum of .3%.  Jeff Karn is
    reviewing the results with Dave Douglas.
QX (Trim Quad)
o  We are integrating the sequestered conductor dimension into our 3D magnetic model and into the design
    drawings.  We are working on a plan to individually match coil size (we have only 8 magnets to build) that
    results from a 140 conductor build up into a variable magnet geometry that will yield the best magnetic
Corrector Dipoles (DB & DJ)
o  The bids were received and are in evaluation.
o  Robin Wines returned from leave to take up design of the sextupole, expanding from her successful 2D design
    into a 3D design in TOSCA.
o  A single Dipole Steel bid was received from a reputable company.  The folks in procurement are following up
    with contacts with the other vendors to get comments.
 o  We continued design layout of the girders between cryomodule.  Many of the interference that we detected in
    the initial layouts were taken care of in a session with Dave Douglas.
o  The Task definition to be used by an engineering service contractor for design of the arc chambers was
    composed and is in checking.  Robby Hicks is putting together the attachments and graphics.
o  We purchased heavy-duty alignment cartridges from CERN (European Organization for Nuclear Research) for
    use under the large 180 degree dipoles.

WBS 10 (Wiggler):
Measurements of the dispersion section were made to characterize the linear behavior of the trim coils.  Data is being reviewed.  We are still waiting for the production trim coils for the wigglers.  Drawings for probe mechanical drive were signed off and parts are out for order.  Still in the works is the drawing for the Hall Probe mount.

WBS 11 (Optics):
The design on the optical cavity assembly internals is proceeding.  During our weekly design review, we made decisions on the mirror fixturing.  For mirror motion, we have drawings on the actuation, but will also pursue a scheme hereby we use a commercially-available motorized actuator.  Several companies produce these, but none will rate their units to a pressure of 10E-8 Torr.  We will borrow one and do base pressure and outgassing tests during the next several weeks.  The Nd:YAG laser for the mirror test stand was tested last Friday.  It made specified power easily, and the interlocks tested OK.  The beampath shields arrived, and one was installed, another is in the shop for modification.  Motion control of the gimbal mount was installed and tested.  Parts to hold our 3" DFM assembly in the 6" gimbal mount were cleaned for HV use and are ready to be used.  We began building a scaled-down version of the mirror metrology system.

The Drive Laser's Nd:YLF laser rod was replaced with our new rod, which will give us better performance.  A bit more alignment work remains to optimize the power.

Chris Behre, an undergraduate student from ODU, joined the Optics Group for the summer.


No operations this week. However, we did have put to put a temporary work procedure in place for working in the FEL Building today and yesterday after a calibration problem was discovered with the oxygen sensors that are used for ODH detectors wherever cryogenic gases are being used.   The calibration problem was noticed during a scheduled spill test in the CEBAF accelerator.  Re-calibrations of the detectors and re-certification of the ODH system are expected by next week.  Many groups in the Accelerator Division worked hard and expeditiously to solve this problem.  Thanks to all involved.